Circular No. 7790

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 Y1 (HELIN-LAWRENCE) On 2001 Dec. 25, S. Nakano (Sumoto, Japan) reported that T. Oribe

had apparently recovered comet P/1993 K2 (= 1993 XI = 1993l) the night before with the 1.0-m reflector at the Saji Observatory. The position was within 2′′ of the prediction by B. G. Marsden on MPC 34423 (ephemeris on MPC 43696). No information was provided about the object’s appearance other than m_{1} = 19.5. The comet has now been independently reported by K. Sárneczky and Z. Heiner in 2002 Jan. 11 data obtained with the 0.6-m Schmidt at Piszkéstető, at m_{1} = 20, but again with no information about the appearance. These observations confirm a tentative single-night detection by C. W. Hergenrother and D. Means of an object of stellar appearance (in an 840-s co-added exposure) at the comet’s expected position a year ago with the Steward Observatory’s 2.3-m reflector at Kitt Peak.

Date UT α_{2000} δ_{2000} m_{2} Observer
2001 Jan. 23.38888 10^{h}38^{m}55.^{s}54 +19^{o}27′58′′.5  Hergenrother
23.39434 10 38 55.39 +19 27 59.6  _{”}
23.39984 10 38 55.32 +19 28 00.1 21.5  _{”}
Dec. 24.85868 13 41 58.34 − 0 41 00.0  Oribe
24.86355 13 41 58.71 − 0 41 02.5  _{”}
24.86946 13 41 58.97 − 0 41 02.8  _{”}
24.87397 13 41 59.23 − 0 41 03.5  _{”}
2002 Jan. 11.19365 13 56 49.99 − 1 30 15.6  Sárneczky
11.19753 13 56 50.23 − 1 30 14.5  _{”}
11.20137 13 56 50.48 − 1 30 16.1  _{”}

Orbital elements by Marsden from 122 observations, 1993–2002 (mean residual 0′′.8):

Epoch = 2003 Jan. 1.0 TT

T = 2002 Dec. 22.4510 TT  ω = 163.^{o}6906
e = 0.307738  Ω = 92.0158 }2000.0
q = 3.110106 AU  i = 9.8713

a = 4.492670 AU n^{o} = 0.1035016 P = 9.523 years

PECULIAR VARIABLE IN MONOCEROS Further to IAUC 7785, J. Bedient adds that the star is not visible on

CCD images in the Stardial archive taken on 141 nights from 1999 Jan. 12.228 to 2001 Dec. 21.287 UT (RG-1 filter, limiting mag ∼ 12).

2002 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green